OpenStudy (sam_ovo):

find all real solutions (x-2)^(4/5)=2

OpenStudy (sam_ovo):

\[(x-2)^{\frac{ 4 }{ 5 }}=2\]

OpenStudy (sshayer):

\[\left\{ \left( x-2 \right)^{\frac{ 4 }{ 5 }} \right\}^{\frac{ 5 }{ 4 }}=2^{\frac{ 5 }{ 4 }}\] can you complete now?

OpenStudy (mathmale):

Hint: \[(x^a)^b = x ^{ab}\]